OpenStudy (anonymous):

my question is: x^+8x+15 / x+5 (polynomial)

OpenStudy (anonymous):

Assuming that the numerator is a polynomial of degree 2: \[\frac{x^{2}+8x+15}{x+5} = x+3\] x + 3 x + 5 | x^2 + 8x + 15 x^2 + 5x _______________ 3x + 15 3x + 15 _________ 0 So you have a quotient of x+3 with 0 remainder.

OpenStudy (anonymous):

Thank you (: